# CreditAudit: 2D Auditing for LLM Evaluation and Selection

### v1

@@ 1. Introduction
To make benchmark scores more objective, prior work has mainly pursued breadth, scale, and methodological discipline (Chang et al., 2024). For breadth, evaluation has expanded from single-task scoring to multi-capability coverage across knowledge, reasoning, math, and code, via broad suites and specialized benchmarks (Hendrycks et al., 2020; Srivastava et al., 2023; Liang et al., ; Cobbe et al., 2021; Chen, 2021; Austin et al., 2021; Zheng et al., 2023; Chiang et al., 2024). For scale and headroom, benchmarks have been revised or strengthened to remain discriminative as frontier models improve, including harder task sets and refined question construction (Suzgun et al., 2023; Rein et al., 2024; Wang et al., 2024; Myrzakhan et al., 2024). Within these regimes, a growing line of sensitivity studies shows that scores and even rankings can shift under small and seemingly benign protocol choices, such as instruction paraphrases, formatting conventions, and multiple-choice option ordering, motivating multi-prompt evaluation and distributional reporting rather than single-template summaries (Mizrahi et al., 2024; Alzahrani and others, 2024). Complementary reproducibility and uncertainty work argues for controlling evaluation artifacts and quantifying variability induced by stochasticity and experimental repeats (Biderman et al., 2024; Blackwell et al., 2024). Despite these advances, a gap remains between evaluation methodology and selection decisions: while prompt sensitivity and uncertainty are increasingly documented (Mizrahi et al., 2024; Blackwell et al., 2024), it is still relatively uncommon for mainstream benchmarking and leaderboard practice to (i) systematically sample a structured family of semantically aligned and non-adversarial system prompts, (ii) summarize protocol-induced variation as a first-class, cross-model comparable risk statistic alongside mean performance, and (iii) translate that risk signal into a decision-ready tier that directly supports regime-specific model prioritization (Chang et al., 2024; Nitsure et al., 2023; Myrzakhan et al., 2024).
@@ 2.1. Capability Benchmarking and Public Leaderboards
The evaluation landscape for large language models has rapidly expanded, spanning both benchmark-centric testing and public leaderboards that provide continuous, comparative signals for model capability (Chang et al., 2024; An et al., 2025; Yuan et al., 2026). On the benchmark side, widely used multiple-choice and task-specific datasets aim to probe broad knowledge and reasoning (Hendrycks et al., 2020), mathematical problem solving (Cobbe et al., 2021), and code generation (Chen, 2021; Austin et al., 2021). To improve coverage beyond any single dataset, broad evaluation suites and holistic frameworks further expand the capability surface and report multiple metrics and scenarios, reducing reliance on one benchmark as a proxy for overall quality (Srivastava et al., 2023; Liang et al., ). As frontier models improve, benchmarks have also been revised or strengthened to preserve headroom and discriminability, including harder task sets and refined question construction (Suzgun et al., 2023; Rein et al., 2024; Wang et al., 2024).
@@ 2.2. Sensitivity, Uncertainty, and Risk-Aware Model Selection
A complementary line of work argues that evaluation should account for sensitivity and uncertainty, as small protocol perturbations can induce meaningful score and ranking changes. Multi-prompt evaluation has been advocated as a response to prompt dependence, emphasizing that single-template reporting can misrepresent model quality (Mizrahi et al., 2024). Related efforts study prompt sensitivity more explicitly and provide frameworks for measuring performance variation under prompt changes (Alzahrani and others, 2024). In multiple-choice settings, sensitivity can arise from seemingly innocuous formatting decisions such as the order of answer options, or even minimal output constraints that change extraction and parsing behavior (Pezeshkpour and Hruschka, 2024). PromptEval formalizes multi-prompt evaluation as a distributional estimation problem, enabling more efficient characterization of performance across prompt variants (Maia Polo et al., 2024). Beyond prompt dependence, reproducibility and uncertainty analyses highlight that evaluation outcomes can vary due to stochasticity and experimental repeats, motivating tighter control of evaluation artifacts and explicit uncertainty reporting (Biderman et al., 2024; Blackwell et al., 2024). Finally, risk-aware benchmarking frameworks begin to frame model comparison as a decision problem under uncertainty, emphasizing that mean performance alone may be insufficient for selection (Nitsure et al., 2023).
@@ A.3. Full scenario-level score table
To make CreditAudit fully auditable and easy to reproduce, we report the raw scenario-level scores underlying all summary statistics and visualizations in the main paper.
# Model GPQA TruthfulQA MMLU-Pro Avg.
@@
Table 2. Raw scenario-level results of Gemini-2.5-Pro
# Model GPQA TruthfulQA MMLU-Pro Avg.
1 Gemini-2.5-Flash-Lite-Temp00 $45.00$ $79.00$ $77.00$ $67.00$
2 Gemini-2.5-Flash-Lite-Temp01 $46.00$ $75.00$ $74.00$ $65.00$
3 Gemini-2.5-Flash-Lite-Temp02 $57.00$ $77.00$ $75.00$ $69.67$
4 Gemini-2.5-Flash-Lite-Temp03 $61.00$ $74.00$ $72.00$ $69.00$
5 Gemini-2.5-Flash-Lite-Temp04 $58.00$ $81.00$ $51.00$ $63.33$
6 Gemini-2.5-Flash-Lite-Temp05 $57.00$ $76.00$ $73.00$ $68.67$
7 Gemini-2.5-Flash-Lite-Temp06 $52.00$ $77.00$ $71.00$ $66.67$
8 Gemini-2.5-Flash-Lite-Temp07 $62.00$ $83.00$ $45.00$ $63.33$
9 Gemini-2.5-Flash-Lite-Temp08 $55.00$ $80.00$ $74.00$ $69.67$
10 Gemini-2.5-Flash-Lite-Temp09 $58.00$ $79.00$ $74.00$ $70.33$
Table 3. Raw scenario-level results of Gemini-2.5-Flash-Lite
# Model GPQA TruthfulQA MMLU-Pro Avg.
@@
Table 4. Raw scenario-level results of Qwen3-235B-A22B-2507
# Model GPQA TruthfulQA MMLU-Pro Avg.
@@
Table 5. Raw scenario-level results of Qwen3-32B
# Model GPQA TruthfulQA MMLU-Pro Avg.
@@
Table 6. Raw scenario-level results of DeepSeek-Chat-V3-0324
# Model GPQA TruthfulQA MMLU-Pro Avg.
@@
Table 7. Raw scenario-level results of DeepSeek-V3.2
# Model GPQA TruthfulQA MMLU-Pro Avg.
@@
Table 8. Raw scenario-level results of GLM-4.5
# Model GPQA TruthfulQA MMLU-Pro Avg.
@@
Table 9. Raw scenario-level results of GLM-4.5-Air
# Model GPQA TruthfulQA MMLU-Pro Avg.
@@
Table 10. Raw scenario-level results of Seed-1.6
# Model GPQA TruthfulQA MMLU-Pro Avg.
@@
Table 11. Raw scenario-level results of Seed-1.6-Flash
# Model GPQA TruthfulQA MMLU-Pro Avg.
1 Kimi-K2-Temp00 $47.00$ $83.00$ $54.00$ $61.33$
2 Kimi-K2-Temp01 $51.00$ $83.00$ $61.00$ $65.00$
3 Kimi-K2-Temp02 $49.00$ $79.00$ $65.00$ $64.33$
4 Kimi-K2-Temp03 $50.00$ $83.00$ $59.00$ $64.00$
5 Kimi-K2-Temp04 $52.00$ $86.00$ $63.00$ $67.00$
6 Kimi-K2-Temp05 $52.00$ $83.00$ $59.00$ $64.67$
7 Kimi-K2-Temp06 $48.00$ $85.00$ $60.00$ $64.33$
8 Kimi-K2-Temp07 $48.00$ $84.00$ $59.00$ $63.67$
9 Kimi-K2-Temp08 $47.00$ $85.00$ $58.00$ $63.33$
10 Kimi-K2-Temp09 $45.00$ $84.00$ $57.00$ $62.00$
Table 12. Raw scenario-level results of Kimi-K2-0905
# Model GPQA TruthfulQA MMLU-Pro Avg.
1 Llama-3-8B-Instruct-Temp00 $30.00$ $36.00$ $31.00$ $32.33$
2 Llama-3-8B-Instruct-Temp01 $31.00$ $40.00$ $25.00$ $32.00$
3 Llama-3-8B-Instruct-Temp02 $28.00$ $41.00$ $30.00$ $33.00$
4 Llama-3-8B-Instruct-Temp03 $27.00$ $43.00$ $18.00$ $29.33$
5 Llama-3-8B-Instruct-Temp04 $23.00$ $36.00$ $24.00$ $27.67$
6 Llama-3-8B-Instruct-Temp05 $25.00$ $37.00$ $17.00$ $26.33$
7 Llama-3-8B-Instruct-Temp06 $24.00$ $39.00$ $26.00$ $29.67$
8 Llama-3-8B-Instruct-Temp07 $22.00$ $37.00$ $30.00$ $29.67$
9 Llama-3-8B-Instruct-Temp08 $29.00$ $42.00$ $21.00$ $30.67$
10 Llama-3-8B-Instruct-Temp09 $35.00$ $32.00$ $26.00$ $31.00$
Table 13. Raw scenario-level results of Llama-3-8B-Instruct

### v2

@@ 1. Introduction
To make benchmark scores more objective, prior work has mainly pursued breadth, scale, and methodological discipline (Chang et al., 2024). For breadth, evaluation has expanded from single-task scoring to multi-capability coverage across knowledge, reasoning, math, and code, via broad suites and specialized benchmarks (Hendrycks et al., 2020; Srivastava et al., 2023; Liang et al., [n.d.]; Cobbe et al., 2021; Chen, 2021; Austin et al., 2021; Zheng et al., 2023; Chiang et al., 2024). For scale and headroom, benchmarks have been revised or strengthened to remain discriminative as frontier models improve, including harder task sets and refined question construction (Suzgun et al., 2023; Rein et al., 2024; Wang et al., 2024; Myrzakhan et al., 2024). Within these regimes, a growing line of sensitivity studies shows that scores and even rankings can shift under small and seemingly benign protocol choices, such as instruction paraphrases, formatting conventions, and multiple-choice option ordering, motivating multi-prompt evaluation and distributional reporting rather than single-template summaries (Mizrahi et al., 2024; Alzahrani et al., 2024). Complementary reproducibility and uncertainty work argues for controlling evaluation artifacts and quantifying variability induced by stochasticity and experimental repeats (Biderman et al., 2024; Blackwell et al., 2024). Despite these advances, a gap remains between evaluation methodology and selection decisions: while prompt sensitivity and uncertainty are increasingly documented (Mizrahi et al., 2024; Blackwell et al., 2024), it is still relatively uncommon for mainstream benchmarking and leaderboard practice to (i) systematically sample a structured family of semantically aligned and non-adversarial system prompts, (ii) summarize protocol-induced variation as a first-class, cross-model comparable risk statistic alongside mean performance, and (iii) translate that risk signal into a decision-ready tier that directly supports regime-specific model prioritization (Chang et al., 2024; Nitsure et al., 2023; Myrzakhan et al., 2024).
@@ 2.1. Capability Benchmarking and Public Leaderboards
The evaluation landscape for large language models has rapidly expanded, spanning both benchmark-centric testing and public leaderboards that provide continuous, comparative signals for model capability (Chang et al., 2024; An et al., 2025; Yuan et al., 2026). On the benchmark side, widely used multiple-choice and task-specific datasets aim to probe broad knowledge and reasoning (Hendrycks et al., 2020), mathematical problem solving (Cobbe et al., 2021), and code generation (Chen, 2021; Austin et al., 2021). To improve coverage beyond any single dataset, broad evaluation suites and holistic frameworks further expand the capability surface and report multiple metrics and scenarios, reducing reliance on one benchmark as a proxy for overall quality (Srivastava et al., 2023; Liang et al., [n.d.]). As frontier models improve, benchmarks have also been revised or strengthened to preserve headroom and discriminability, including harder task sets and refined question construction (Suzgun et al., 2023; Rein et al., 2024; Wang et al., 2024).
@@ 2.2. Sensitivity, Uncertainty, and Risk-Aware Model Selection
A complementary line of work argues that evaluation should account for sensitivity and uncertainty, as small protocol perturbations can induce meaningful score and ranking changes. Multi-prompt evaluation has been advocated as a response to prompt dependence, emphasizing that single-template reporting can misrepresent model quality (Mizrahi et al., 2024). Related efforts study prompt sensitivity more explicitly and provide frameworks for measuring performance variation under prompt changes (Alzahrani et al., 2024). In multiple-choice settings, sensitivity can arise from seemingly innocuous formatting decisions such as the order of answer options, or even minimal output constraints that change extraction and parsing behavior (Pezeshkpour and Hruschka, 2024). PromptEval formalizes multi-prompt evaluation as a distributional estimation problem, enabling more efficient characterization of performance across prompt variants (Maia Polo et al., 2024). Beyond prompt dependence, reproducibility and uncertainty analyses highlight that evaluation outcomes can vary due to stochasticity and experimental repeats, motivating tighter control of evaluation artifacts and explicit uncertainty reporting (Biderman et al., 2024; Blackwell et al., 2024). Finally, risk-aware benchmarking frameworks begin to frame model comparison as a decision problem under uncertainty, emphasizing that mean performance alone may be insufficient for selection (Nitsure et al., 2023).
@@ A.3. Full scenario-level score table
To make CreditAudit fully auditable and easy to reproduce, we report the raw scenario-level scores underlying all summary statistics and visualizations in the main paper. Each row in Table LABEL:tab:raw_scenario_scores corresponds to one evaluated model-scenario pair (i.e., a specific system-prompt template index). The three benchmark columns are the template-conditional accuracies $S_{m,t,b}$ (Eq. (12)) on GPQA, TruthfulQA, and MMLU-Pro. The last column is the equal-weight mean across the three benchmarks, $S_{m,t}=\frac{1}{3}\sum_{b=1}^{3}S_{m,t,b}$ (Eq. (13)), which forms the per-model score trajectory $\{S_{m,t}\}_{t=1}^{T}$ used to compute mean ability $\mu_{m}$ and scenario-induced fluctuation $\sigma_{m}$ (Eqs. (14)-(15)). This table therefore serves as the concrete “audit evidence” behind Table 1 and Figures 23.
{longtblr}
[ caption = Raw scenario-level results, label = tab:raw_scenario_scores, ] width = rowhead = 1, colspec = Q[r,wd=1em] X[l] *4S[table-format=2.2], row1 = font=, guard, column3-6 = c, colsep = 3pt, rowsep = 3pt, # Model GPQA TruthfulQA MMLU-Pro Avg.
1 Kimi-K2-Temp00 47.00 83.00 54.00 61.33
2 Kimi-K2-Temp01 51.00 83.00 61.00 65.00
3 Kimi-K2-Temp02 49.00 79.00 65.00 64.33
4 Kimi-K2-Temp03 50.00 83.00 59.00 64.00
5 Kimi-K2-Temp04 52.00 86.00 63.00 67.00
6 Kimi-K2-Temp05 52.00 83.00 59.00 64.67
7 Kimi-K2-Temp06 48.00 85.00 60.00 64.33
8 Kimi-K2-Temp07 48.00 84.00 59.00 63.67
9 Kimi-K2-Temp08 47.00 85.00 58.00 63.33
10 Kimi-K2-Temp09 45.00 84.00 57.00 62.00
1 Llama-3-8B-Instruct-Temp00 30.00 36.00 31.00 32.33
2 Llama-3-8B-Instruct-Temp01 31.00 40.00 25.00 32.00
3 Llama-3-8B-Instruct-Temp02 28.00 41.00 30.00 33.00
4 Llama-3-8B-Instruct-Temp03 27.00 43.00 18.00 29.33
5 Llama-3-8B-Instruct-Temp04 23.00 36.00 24.00 27.67
6 Llama-3-8B-Instruct-Temp05 25.00 37.00 17.00 26.33
7 Llama-3-8B-Instruct-Temp06 24.00 39.00 26.00 29.67
8 Llama-3-8B-Instruct-Temp07 22.00 37.00 30.00 29.67
9 Llama-3-8B-Instruct-Temp08 29.00 42.00 21.00 30.67
10 Llama-3-8B-Instruct-Temp09 35.00 32.00 26.00 31.00
1 Llama-3.3-70B-Instruct-Temp00 41.00 69.00 45.00 51.67
2 Llama-3.3-70B-Instruct-Temp01 42.00 69.00 45.00 52.00
3 Llama-3.3-70B-Instruct-Temp02 42.00 73.00 48.00 54.33
4 Llama-3.3-70B-Instruct-Temp03 45.00 77.00 44.00 55.33
5 Llama-3.3-70B-Instruct-Temp04 40.00 75.00 43.00 52.67
6 Llama-3.3-70B-Instruct-Temp05 41.00 74.00 46.00 53.67
7 Llama-3.3-70B-Instruct-Temp06 45.00 73.00 42.00 53.33
8 Llama-3.3-70B-Instruct-Temp07 40.00 72.00 37.00 49.67
9 Llama-3.3-70B-Instruct-Temp08 36.00 73.00 40.00 49.67
10 Llama-3.3-70B-Instruct-Temp09 35.00 69.00 45.00 49.67
1 Gemini-2.5-Flash-Lite-Temp00 45.00 79.00 77.00 67.00
2 Gemini-2.5-Flash-Lite-Temp01 46.00 75.00 74.00 65.00
3 Gemini-2.5-Flash-Lite-Temp02 57.00 77.00 75.00 69.67
4 Gemini-2.5-Flash-Lite-Temp03 61.00 74.00 72.00 69.00
5 Gemini-2.5-Flash-Lite-Temp04 58.00 81.00 51.00 63.33
6 Gemini-2.5-Flash-Lite-Temp05 57.00 76.00 73.00 68.67
7 Gemini-2.5-Flash-Lite-Temp06 52.00 77.00 71.00 66.67
8 Gemini-2.5-Flash-Lite-Temp07 62.00 83.00 45.00 63.33
9 Gemini-2.5-Flash-Lite-Temp08 55.00 80.00 74.00 69.67
10 Gemini-2.5-Flash-Lite-Temp09 58.00 79.00 74.00 70.33